TWO JETS LEAVE DENVER AT 9 AM ONE FLYING EAST AT A SPEED 50km/h GREATER THAN THE OTHER,WHICH IS TRAVELING WEST.AT 11 AM THE PLANES ARE 2500 km APART .FIND THERE SPEEDS.
I HAVE TRIED SEVERAL WAYS IS THERE AN EASIER WAY TO REMEMBER THE FORMULAS FOR THIS

Answer by stanbon(48558)

TWO JETS LEAVE DENVER AT 9 AM ONE FLYING EAST AT A SPEED 50km/h GREATER THAN THE OTHER,WHICH IS TRAVELING WEST.AT 11 AM THE PLANES ARE 2500 km APART .FIND THERE SPEEDS.
---------------------
East jet DATA:
rate = (x+50) km/h ; time = 2 hr ; distance = rt = 2(x+50) km
-------------------------------
West jet DATA:
rate = x km/h ; time = 2 hr ; distanc e= 2x km
-----------------------------
Equation:
distance + distance = 2500
2(x+50) + 2x = 2500
4x = 2400
x = 600 km/h (speed of the west jet)
x+50 = 650 km/h (speed of the east jet)
===========================================
Cheers,
Stan H.